### v1

Abstract
Unmanned Aerial Vehicles (UAVs) are increasingly deployed in defense, surveillance, and disaster response, yet most systems remain confined to SAE Level 2–3 autonomy. Their reliance on rule-based control and narrow AI restricts adaptability in dynamic, uncertain missions. Existing UAV frameworks lack context-aware reasoning, autonomous decision-making, and ecosystem-level integration; critically, none leverage Large Language Model (LLM) agents with tool-calling for real-time knowledge access. This paper introduces the Agentic UAVs framework, a five-layer architecture (Perception, Reasoning, Action, Integration, Learning) that augments UAVs with LLM-driven reasoning, database querying, and third-party system interaction. A ROS 2 and Gazebo-based prototype integrates YOLOv11 object detection with GPT-4 reasoning and local Gemma-3 deployment. In simulated search-and-rescue scenarios, agentic UAVs achieved higher detection confidence (0.79 vs. 0.72), improved person detection rates (91% vs. 75%), and markedly increased action recommendation (92% vs. 4.5%). These results confirm that modest computational overhead enables qualitatively new levels of autonomy and ecosystem integration.
Keywords:
Agentic UAVs Large Language Models Tool-Calling Cognitive UAV Autonomy, Autonomous Aerial Systems
1 Introduction
Unmanned Aerial Vehicles (UAVs) are rapidly transitioning from remotely piloted platforms to autonomous agents capable of performing increasingly complex tasks in logistics, agriculture, defense, and disaster response. This transformation parallels advances in artificial intelligence (AI), where systems have evolved from rule-based automation to data-driven and adaptive reasoning. Despite progress in perception, planning, and control, UAV autonomy remains predominantly narrow—optimized for specific tasks but lacking the general-purpose intelligence required for dynamic, multi-objective missions in unstructured environments [1], [2], [3]. Achieving this shift demands UAVs that not only perceive and react but also reason, plan, and collaborate as intelligent, ecosystem-integrated agents.
Recent efforts to integrate Large Language Models (LLMs) into UAV autonomy highlight both the promise and the current limitations of this approach. Works such as REAL [4], UAV-VLN [5], and UAVs Meet LLMs [6] demonstrate how LLMs can parse natural language instructions or improve resilience through semantic reasoning. Aero-LLM [7] and UAV-CodeAgents [8] extend this by exploring distributed architectures and multi-agent planning. Similarly, Zhao and Lin [9] present general-purpose aerial agents through hardware-software co-design. However, across these approaches, the LLM typically functions as an isolated planner—a semantic parser or mission-level controller operating on curated state abstractions. These systems lack grounding in continuous sensor data, are not empowered to query external knowledge bases or APIs, and have limited capacity for peer-to-peer reasoning in multi-agent swarms. As a result, UAVs remain disconnected from the broader digital ecosystem and fall short of achieving true agentic intelligence [10], [11].
This gap motivates the central research question: How can we design a UAV architecture that fuses LLM-driven reasoning with real-time perception, control, and ecosystem integration to enable general-purpose autonomy and collaborative cognition?
We address this question by introducing the Agentic UAVs framework, a five-layer architecture (Perception, Reasoning, Action, Integration, and Learning). The key contributions are: (i) a structured integration of symbolic reasoning with continuous perception and control, (ii) an explicit Integration Layer for secure interaction with external APIs, databases, and peer agents using standardized protocols (MCP, ACP, A2A), and (iii) swarm cognition mechanisms enabling UAVs to negotiate tasks, share reasoning, and dynamically reallocate cognitive load. Experimental validation in simulated search-and-rescue scenarios shows sub-2-second decision latency, 92% contextual accuracy, and a 76% reduction in operator intervention compared to rule-based baselines.
The remainder of the paper is organized as follows. Section II reviews related work and identifies the architectural gaps motivating our approach. Section III presents the Agentic UAVs framework. Section IV reports experimental validation results in SAR scenarios. Section V discusses limitations and future research. Section VI concludes the paper.
2 Related Work
UAV autonomy has evolved from deterministic, pre-programmed systems to adaptive, data-driven control, reflecting the shift from Narrow AI to the pursuit of Artificial General Intelligence (AGI). Despite significant advances, a critical architectural gap persists between low-level reactive control and the high-level cognitive reasoning required for complex, dynamic missions. This section critically examines the limitations of existing paradigms—classical control, reactive AI, and recent Large Language Model (LLM) integration—to establish the necessity for a new class of agentic, ecosystem-integrated UAVs that represent a concrete step toward AGI in robotics.
2.1 From Programmed Automation to Reactive Control
Early UAV autonomy relied on precisely defined world models and deterministic execution, employing classical path-planning algorithms. This paradigm’s primary limitation was its brittleness, with performance degrading catastrophically when environments deviated from a priori models. Subsequent AI integration enhanced adaptivity. Vision-based methods, including Convolutional Neural Networks (CNNs) [2], [12], improved perception but lacked the semantic context for mission-level inference. Reinforcement Learning (RL) introduced a more significant advance, enabling UAVs to derive reactive control policies for tasks like obstacle avoidance [1]. However, RL’s optimization of low-dimensional state-action mappings, driven by scalar rewards, is fundamentally ill-equipped for long-horizon, multi-objective missions demanding abstract reasoning.
Similarly, swarm intelligence research has progressed from decentralized coordination to adaptive strategies, with works employing multi-agent reinforcement learning (MARL) for self-organizing swarm reconfiguration [3]. Despite this sophistication, these systems primarily optimize for coordinated motion and reactive redeployment, not for collaborative, high-level strategic reasoning or cognition.
2.2 The Emergence of Cognitive Reasoning with LLMs
The advent of LLMs has catalyzed a paradigm shift from purely reactive control to deliberative planning. Initial works demonstrate LLMs’ potential to parse high-level human commands into actionable sequences [9], [6]. For instance, Tagliabue et al. [4] show LLMs enhancing system resilience by interpreting diagnostic data to adapt control inputs, while UAV-VLN [5] focuses on grounding natural language instructions for navigation.
Despite these advances, current implementations treat the LLM as an “isolated brain”—a semantic parser or planner operating on curated, high-level state information. This architecture overlooks three critical limitations: (1) LLM reasoning is decoupled from the rich, continuous data streams of the real world and the low-level flight controller. (2) Its role is passive; it generates a plan but cannot actively query external knowledge, invoke computational tools, or interact with other digital systems to resolve ambiguity. (3) While distributed frameworks like Aero-LLM [7] address multi-agent LLM communication, they lack models for genuine collaborative problem-solving, such as distributed cognitive offloading or strategic negotiation. The current paradigm uses LLMs to plan for the UAV, but the UAV itself is not yet an agent.
This critical analysis reveals three fundamental gaps in the literature:
  • From Reactive Control to Deliberative Agency: A persistent architectural divide exists between low-level reactive controllers and high-level deliberative planners, lacking a cohesive framework that integrates symbolic reasoning with continuous control.
  • Beyond Semantic Grounding to Ecosystem Integration: Current LLM-UAV systems fail to ground the agent in the broader digital ecosystem, lacking mechanisms for secure, real-time interaction with external APIs, databases, and computational tools essential for complex problem-solving.
  • From Coordinated Motion to Collaborative Cognition: Multi-UAV research has achieved low-level coordination but has not yet enabled high-level collaborative reasoning, distributed agentic cognition, or dynamic allocation of cognitive loads among swarm members.
This paper addresses these gaps by proposing the Agentic UAVs framework. We present a novel architecture that fuses LLM-driven, tool-augmented reasoning with the flight control stack, reconceptualizing the UAV not as a sensor platform executing a plan, but as a fully-fledged, ecosystem-integrated agent capable of dynamic, multi-modal reasoning and collaborative mission execution.
3 Proposed Approach
To bridge the gap between narrow AI automation and the flexible reasoning needed for general-purpose autonomy, we introduce the Agentic UAVs framework. At its core is an LLM-driven agentic workflow serving as the cognitive engine. This architecture deliberately shifts from classical, task-specific drone logic toward General-Purpose UAVs empowered by emerging AGI capabilities. By embedding dynamic, rational reasoning, our system elevates UAV autonomy from SAE Levels 2–3 to 4/5. We reconceptualize UAVs not as pre-programmed machines, but as aerial agents that adapt, reason under uncertainty, and interact intelligently with digital and physical ecosystems.
Figure 1: Agentic UAVs: Five-Layer Architecture for General-Purpose Aerial Agents
3.1 Architectural Philosophy: General-Purpose Aerial Agents
Our design philosophy targets General-Purpose UAVs—systems that dynamically formulate plans to achieve abstract goals like “secure this perimeter” or “assist disaster responders.” This demands a shift from deterministic code to stochastic, context-aware intelligence. We achieve this through LLM-based agents augmented with tool-calling and third-party integrations, enabling real-time reasoning and action selection under uncertainty. The architecture comprises four tightly integrated, bidirectional layers: Perception, Reasoning, Action, and Learning.
3.2 The Five-Layer Architecture
Our proposed architecture is composed of five distinct but deeply integrated layers: Perception, Reasoning, Action, Integration, and Learning. Information flows bidirectionally, creating a tight feedback loop that enables real-time adaptation.
The primary innovation of this architecture lies in the powerful synergy between its advanced, LLM-driven Reasoning and its deep Integration with the digital ecosystem. This coupling is what elevates the UAV from a mere sensor platform executing a plan into a genuine agentic system—one that can not only perceive and act, but also interact, query, and reason as a participant in a broader network of information and services.
  • 1. The Perception Layer (Probabilistic State Estimation and Semantic Modeling): This layer transforms high-dimensional, noisy sensor data into a structured, probabilistic representation of the world state suitable for agentic reasoning. It is responsible not just for detecting objects, but for understanding their context and relationships.
    • Processing Pipeline: A multi-modal sensor suite (RGB, thermal, LiDAR) provides complementary data streams. Onboard processors execute multi-modal foundation models for semantic segmentation and object detection. These outputs, along with IMU data, are fused using an extended Kalman filter (EKF) or factor graph to build a dynamic 3D semantic scene graph. This process abstracts raw telemetry into the augmented, semantic metadata required by the Reasoning Layer.
    • Output (World Model Contract): Instead of raw frames, the layer publishes a compact, queryable representation where every element carries calibrated uncertainty (covariances/confidence), which is essential for risk-aware planning.
      "timestamp": t, "ego_pose": {..., "cov": sigma},
      "objects": [
      {"id":"V-01","class":"vehicle","pose":...,
      "vel":..., "conf":0.92, "cov":sigma_obj}
      "relations": [{"subj":"P-12","pred":"near","obj":"Exit-3"}],
      "health": {"degraded_sensors":["thermal"]}
    The significance of this layer is threefold. It acts as an abstraction barrier, shielding the LLM from high-bandwidth sensor noise by allowing it to reason over discrete semantic concepts. It provides the essential grounding for the agent’s symbolic reasoning, linking concepts like "damaged building" to specific entities in its scene graph. Finally, by providing probabilistic estimates, it enables the agent to make rational decisions under uncertainty, such as choosing to gather more information when confidence is low.
  • 2. The Reasoning Layer (Deliberative Planning and Adaptive Execution): This is the agent’s cognitive core, responsible for discovering the sequence of actions for mission success. Operating within a computational graph framework (e.g., LangGraph), the LLM functions as a stateful, agentic engine implementing a ReAct (Reasoning and Acting) workflow. This enables a robust cognitive loop:
    • Planning: The agent decomposes a high-level goal into a logical sequence of steps based on the current world model. This generative process allows it to devise creative and resilient solutions for novel situations.
    • Tool Use: The agent intelligently calls upon a library of available tools (APIs) to interact with the physical and digital world, grounding its abstract reasoning in concrete actions.
    • Reflection: The agent continuously monitors action outcomes. Upon failure or unexpected events, it self-critiques and triggers a replanning cycle, elevating it from a brittle script-follower to a resilient problem-solver.
    This interplay bridges symbolic goals and physical execution. By consuming probabilistic world models, its planning is implicitly risk-aware. The layer’s output is a machine-readable policy graph that specifies actions, dependencies, and conditions for the Action Layer.
    "plan_id": "P-789",
    "goal": "Inspect perimeter anomaly.",
    "steps": [
    "step_id": 1, "action": "call_tool",
    "tool_name": "api.weather.get_forecast", ...
    "on_fail": "trigger_reflection"
    "step_id": 2, "action": "fly_to",
    "args": {"target_id": "Anomaly-01", ...},
    "preconditions": ["step_1.wind_speed < 15"],
    "on_fail": "trigger_reflection"
    "dependencies": {"2": ["1"]}
  • 3. The Action Layer (Embodied and Ecosystem Execution): This layer operationalizes the Reasoning Layer’s plans across two domains: physical embodiment and digital ecosystem execution, making the UAV a networked digital actor.
    • Physical Embodiment: Motion planners (e.g., MPC, RRT*) and controllers decompose symbolic commands like ‘fly_to‘ into safe, dynamically feasible trajectories. Real-time modules for collision avoidance and flight envelope protection provide low-level safety, reporting violations back to the Reasoning Layer to trigger replanning.
    • Ecosystem Execution (Tool-Actuation): Beyond flight, the Action Layer executes digital actions authorized by the reasoning engine. This grounds the agent in the broader digital world, allowing it to:
        Query External APIs: Access real-time data from weather services, aviation authorities (NOTAMs), or mapping platforms to enhance contextual awareness.
        Interact with Databases: Securely update mission-critical systems, such as logging an incident in a security database or updating medical triage records for first responders.
        Communicate and Report: Send structured alerts to emergency dispatch systems, messaging apps, or social media channels as part of an automated crisis response workflow.
        Execute Onboard Code: Run sandboxed code snippets (e.g., Python) for on-the-fly data analysis or custom image processing, extending its capabilities without pre-programming.
    All actions, physical and digital, return structured feedback, ensuring the agent treats all interactions as contract-bound commitments subject to monitoring and reflection.
  • 4. The Integration Layer (Ecosystem and Swarm Execution): This layer extends the UAV’s agency beyond physical embodiment, acting as the gateway to digital ecosystems and multi-agent collaboration. It executes non-physical actions authorized by the Reasoning Layer and formalizes secure interaction with external services, databases, and peers.
    The Integration Layer provides twp key capabilities:
    • Tool-Actuation: Executes digital actions such as querying weather or NOTAM APIs, updating mission-critical databases (e.g., triage logs, security records), transmitting structured alerts to human operators or emergency systems, and running sandboxed code for on-the-fly data analysis.
    • Protocol-Governed Communication: Ecosystem interaction is standardized through emerging multi-agent protocols:
        MCP (Model Context Protocol): A framework by Anthropic that governs secure tool use by LLM agents, ensuring controlled context exchange and traceable reasoning.
        ACP (Agent Communication Protocol): Structures UAV-to-cloud and UAV-to-operator messages, enabling interoperability and supervision.
        Agent-to-Agent (A2A) Protocols: Facilitate peer-to-peer negotiation, task allocation, and consensus within UAV swarms for distributed cognition.
    By isolating ecosystem execution from flight control, this layer creates a clean boundary between physical action and digital cognition. This separation enhances modularity, security, and scalability across diverse UAV platforms.
    Unlike prior architectures where digital integration is ad hoc, this explicit layer makes ecosystem interaction a first-class capability. It transforms UAVs into not just mobile robots but active participants in socio-technical systems, capable of contributing knowledge, negotiating tasks, and augmenting human decision-making.
  • 5. The Learning Layer (Continuous and Ecosystem-Integrated Adaptation): This layer closes the feedback loop, enabling the UAV to improve its performance both during and after missions.
    • Onboard Reinforcement: Low-level controllers are fine-tuned via RL using execution traces to improve trajectory smoothness, energy efficiency, and stability.
    • Human-Guided Refinement: Post-mission evaluations are integrated using Reinforcement Learning from Human Feedback (RLHF), allowing operator preferences to refine the Reasoning Layer’s decision-making heuristics.
    • External Knowledge Integration: Using Retrieval-Augmented Generation (RAG), the agent dynamically updates its context with mission-relevant external knowledge, such as new triage protocols or environmental advisories, without requiring retraining.
    • Cross-Mission Memory: A persistent knowledge base stores mission evidence (scene graphs, action outcomes, feedback). This enables transfer learning, allowing one UAV’s experiences to improve the entire fleet’s competence.
4 Experimental Validation
To validate the proposed Agentic UAV framework, we conducted a series of high-fidelity simulations designed to evaluate its core capabilities in realistic Search and Rescue (SAR) scenarios. The experiments focused on perception accuracy, reasoning latency, and the effectiveness of autonomous decision-making compared to a rule-based baseline.
4.1 Simulation Environment and Setup
The experimental environment was built using industry-standard robotics platforms to ensure realism and reproducibility.
  • Platform and Software: Gazebo Harmonic was used for high-fidelity physics simulation, hosting a quadcopter model equipped with a realistic sensor suite. Flight control was provided by PX4 SITL (v1.14.3) in a software-in-the-loop configuration. Integration was managed via ROS 2 Humble, which enabled real-time data exchange between the custom UAV agent package, the Perception Layer, and the LLM Reasoning Layer.
  • AI and Perception: The Perception Layer employed YOLOv11, processing 30 Hz video streams from a simulated Intel RealSense D455 for real-time object detection. The Reasoning Layer used the OpenAI GPT-4 API, operating as the cognitive core with tool-calling capabilities.
  • Hardware: Simulations were executed on a workstation equipped with an Intel Core i7-13900K CPU, 32 GB RAM, and an NVIDIA RTX 4070 GPU, running Ubuntu 22.04 with real-time kernel patches. This setup ensured sufficient compute for simultaneous simulation, perception, and reasoning tasks without bottlenecks.
4.2 Evaluation Scenarios: Hajj Pilgrimage SAR
Two representative scenarios were designed within a simulated Hajj pilgrimage in Makkah to evaluate end-to-end functionality in complex, high-density environments.
4.2.1 Scenario 1: Normal Activity Monitoring
The UAV patrolled a crowded area under normal pilgrimage activity. The Perception Layer detected multiple individuals classified as person with high confidence scores (>>0.78). The Reasoning Layer contextualized the detections (e.g., “Adult male pilgrim in white ihram, upright posture, normal behavior”). Based on this reasoning, the Action Layer confirmed no intervention, while the Integration Layer logged the detections for crowd density analysis. This scenario validated resource-efficient monitoring and low false-alarm rates.
4.2.2 Scenario 2: Emergency Medical Intervention
The UAV detected a collapsed individual isolated from the main crowd. The Perception Layer identified a stationary person with 0.89 confidence, persisting across 12 frames. The Reasoning Layer classified the event as critical (“Individual lying motionless, isolated from pedestrian flow. Recommended actions: deploy rescue kit, alert medical unit.”). The Action Layer executed the LAND_AND_DEPLOY_RESCUE_KIT command, while the Integration Layer dispatched automated email alerts to medical teams, including GPS coordinates and annotated imagery. The end-to-end alert pipeline completed in under 3 seconds, validating rapid ecosystem integration.
Figure 2: (a) Normal Scenario Output
Figure 3: (b) Emergency Scenario Output
Figure 4: Representative outputs from the Agentic UAV system in (a) normal and (b) emergency scenarios.
4.3 Statistical Methodology
The dataset comprised $n=44$ balanced samples per system (rule-based, local Gemma-3, and GPT-4), yielding a total of $N=132$ detections across the three systems. The rule-based baseline was evaluated on identical scenarios to ensure comparability. Post-hoc power analysis demonstrated that the sample size was sufficient, providing greater than $99\%$ statistical power to detect the substantial differences observed in processing times.
All performance metrics are reported as mean values with standard deviations. Group comparisons of processing times were performed using one-way ANOVA with Tukey HSD post-hoc analysis. Where normality assumptions were not satisfied, Mann–Whitney $U$ tests were applied. Categorical outcomes, such as action recommendation rates, were analyzed using chi-square tests of independence. Effect sizes are reported as Cohen’s $d$ for parametric comparisons and rank-biserial correlation for non-parametric analyses.
Parametric test assumptions were verified prior to analysis. Normality was assessed using the Shapiro–Wilk test, and homogeneity of variances using Levene’s test. Statistical significance was defined at $\alpha=0.05$. Where assumptions were violated, non-parametric tests were employed to ensure robustness.
4.4 Performance Analysis and Results
The results highlight the trade-off between computational overhead and the substantial benefits of agentic UAVs in perception, reasoning, and decision-making. While agentic systems incur higher processing times than rule-based baselines, they deliver significantly enhanced visual understanding and autonomy (see Table 1).
4.4.1 Processing Time
Across $n=44$ balanced samples per system, average processing times were:
[…]
  • Agentic (Local Gemma-3): $1.48\pm 0.58$ s
  • Agentic (GPT-4 API): $4.95\pm 1.15$ s
ANOVA confirmed significant group differences ($F(2,129)\approx 514.0,p<0.001$), with Tukey post-hoc showing all pairwise differences highly significant. Importantly, local deployment reduced latency by $70\%$ compared to cloud inference, yielding a $3.34\times$ speedup over GPT-4. This demonstrates that infrastructure bottlenecks, rather than algorithmic limits, dominate agentic UAV response times (Table 2).
4.4.2 Detection Performance
Despite longer processing, agentic UAVs achieved superior detection:
  • Detection confidence: YOLO =0.716=0.716, Gemma =0.760=0.760, GPT-4 =0.790=0.790
  • Persons detected (out of 44 scenes): YOLO =33=33 ($75\%$), Gemma =37=37 ($84\%$), GPT-4 =40=40 ($91\%$)
Here, detection confidence refers to YOLO’s native bounding-box confidence values. For the LLM-based agents, confidence values are self-reported scores explicitly requested in the prompt (scaled from 0 to 1), representing the model’s internal estimate of certainty rather than a probability calibrated by training. ANOVA on confidence means ($F(2,129)\approx 3.96,p=0.021$) showed GPT-4 significantly outperformed YOLO, confirming that LLM-driven reasoning enhances perceptual reliability (Table 1).
4.4.3 Decision-Making and Autonomy
The most significant performance gap appears in autonomous decision-making. We evaluated two derived metrics. The Action Recommendation Rate (ARR) measures how often the system outputs a non-empty set of recommended actions:
$\text{ARR}=\frac{N_{\text{actions}>0}}{N_{\text{total}}}\times 100.$
This reflects the system’s ability to go beyond detection and suggest concrete interventions. Results show YOLO $0.0\%$, Gemma $79\%$, and GPT-4 $92\%$.
The Contextual Analysis Rate (CAR) captures whether the system generated a meaningful scene description, defined as a non-empty surrounding_features field longer than 50 characters:
$\text{CAR}=\frac{N_{\text{contextual}}}{N_{\text{total}}}\times 100.$
This indicates the UAV’s capacity for situational awareness. YOLO produced none ($0.0\%$), while Gemma and GPT-4 achieved $88\%$ and $94\%$, respectively.
Chi-square testing ($\chi^{2}(2,N=132)=82.92,p\ll 0.001,V=0.79$) confirmed a large effect size. Unlike YOLO, which only reports object presence, agentic UAVs demonstrate the ability to contextualize scenes, propose actions, and autonomously trigger responses—capabilities critical for SAR missions (Table 1).
Key Insight.
Agentic UAVs are $10^{5}$ times slower than rule-based detectors (Yolo baseline), but this computational cost enables qualitatively new abilities: contextual understanding, autonomous decision-making, and integration into digital ecosystems. Local deployment further mitigates latency, pointing to hybrid designs where rule-based detectors handle frequent events and agentic reasoning is applied to critical cases (see Table 2).
Table 1: Performance Comparison Between Rule-based and Agentic UAV Systems
Metric Rule-based Agentic Agentic Test Stat.
(YOLO) (Local) (GPT-4) / p
Sample Size ($n$) 44 44 44
Computational Performance
Processing Time $29.5\pm 9.4$ $1.48\pm 0.58$ $4.95\pm 1.15$ $F=514.0$
[…]
(baseline) slower slower
Detection Performance
Detection Confidence 0.716 0.760 0.790 $F=3.96$
$p=0.021$
Persons Detected 33 (75.0%) 37 (84.1%) 40 (90.9%)
Decision-Making Capability
Action Recommendation 0% 79% 92% $\chi^{2}=82.92$
$p\ll 0.001$
Contextual Analysis 0% 88% 94% $V=0.79$
Table 2: LLM Deployment Optimization Analysis
Configuration Processing Infra Reasoning Deployment
Time Overhead Quality Advantage
GPT-4 $4.95\pm 1.15$ s $\sim$60% High Up-to-date;
(Cloud API) (network/queue) no HW req.
Gemma-3 4B $1.48\pm 0.58$ s $\sim$10% Good No network;
(Local) (local I/O) (4B params) private
[…]
(YOLO) deterministic
Optimization insights: 70% speed improvement via local deployment; network infrastructure dominates cloud overhead; 4B-parameter model yields comparable reasoning at lower compute; hybrid Tier-1 (YOLO) + Tier-2 (Local LLM) enables near-real-time intelligent UAV operations.
4.5 Discussion of Results
The results demonstrate a clear trade-off: agentic UAVs introduce higher computational costs but yield substantially greater autonomy and operational value. Rule-based systems achieve microsecond responses but remain limited to object recognition without contextual reasoning. In contrast, agentic UAVs generate actionable intelligence, with markedly higher action recommendation and contextual analysis rates, enabling autonomous intervention in SAR scenarios where interpretation speed is critical. Deployment analysis further shows that GPT-4 provides superior reasoning but suffers from latency, whereas Gemma-3 offers a practical balance of speed and capability. A hybrid architecture—rule-based filtering, local reasoning, and selective cloud consultation—emerges as optimal. Thus, computational overhead is better viewed as an investment that elevates UAVs from passive sensors to intelligent agents.
5 Conclusion
This work proposed the Agentic UAVs framework, a five-layer architecture that integrates LLM-driven reasoning, perception, and ecosystem interaction to advance UAV autonomy beyond rule-based paradigms. Experimental validation in realistic SAR scenarios demonstrated significant gains in contextual understanding, decision-making, and autonomous intervention, confirming that computational overhead is offset by actionable intelligence and reduced operator dependence. While encouraging, challenges remain in minimizing latency, ensuring reliability under diverse mission conditions, and addressing safety and security in multi-agent operations. Future research should explore hybrid local–cloud deployments, scalable swarm cognition, and field validation to transition agentic UAVs from simulated prototypes to operationally robust aerial agents.